OpenStudy (anonymous):

sin2x=cosx on [0,2pi]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

sin 2x = cos x sin 2x - cos x = 0 2 sin x cos x - cos x = 0 cos x (2 sin x - 1) = 0 cos x = 0 or sin x = 1/2 Between 0 and 2 pi : x = pi/2 ; 3pi/2 ; pi/6 ; 5pi/6 I hope this is what you are looking for.

OpenStudy (anonymous):

Thank you!! i got it xD

OpenStudy (anonymous):

You're Welcome !